OpenStudy (moongazer):

A 250-N crate is pulled along a horizontal sidewalk at a constant speed by a rope which makes an angle of 40 degrees with the sidewalk. If a force is 150-N is applied on the rope what is the coefficient of sliding friction?

OpenStudy (anonymous):

μmg=fcosθ μ=fcos40/mg u = 150\cos40/250 \] μ=(3/5)cos40=0.459

OpenStudy (moongazer):

please answer my last question there :)